OpenStudy (anonymous):

What is the sum of the first 35 numbers in the arithmetic series 124 + 133 + 142 + 151 + ⋯? A. 5,690 B. 9,566 C. 9,659 D. 9,695 E. 10,010

OpenStudy (anonymous):

a=124 , d=9 (the difference) , n = 35 plug them in the formula: S=n/2(2a + (n-1)d)

OpenStudy (anonymous):

It is D

OpenStudy (anonymous):

Check it for yourself

OpenStudy (anonymous):

a1 = 124 d = 9 n = 35 an = a1 + (n - 1) * d a35 = 124 + (35 - 1) * 9 a35 = 124 + 34 * 9 a35 = 124 + 306 a35 = 430 (this is just the 35 term) sn = n(a1 + an) / 2 s35 = 35(124 + 430) / 2 s35 = 35 (554)/2 s35 = 19390/2 s35 = 9695 I agree....answer is D